SOLVING THE MYSTERY OF GREEN BANDS OF YTTRIUM MONOHYDRIDE WITH THE HELP OF STARK SPECTROSCOPY.

\underline{ZYGMUNT J. JAKUBEK} and BENOIT SIMARD, Steacie Institute for Molecular Sciences, National Research Council, Ottawa, ON, Canada K1A 0R6; WALTER J. BALFOUR, Department of Chemistry, University of Victoria, Victoria, BC, Canada V8W 3V6.

Stark spectra for four bands of yttrium hydride in the 19300 − 19800 cm^{−1} region have been recorded by exciting the molecules in the presence of external electric field up to 16 kV/cm and monitoring laser induced fluorescence. The molecules have been produced in a free jet molecular beam apparatus by laser vaporizing yttrium metal in the presence of He carrier gas doped with 2 − 5% of H_{2} or NH_{3}. Rotational and Ω quantum number assignments of the bands have been established with the help of Stark effect. The electric dipole moments have been measured to be equal to: 1.569(64), 2.325(8), 3.316(11), 2.273(6), and 1.537(8) for the X^{1}Σ^{+} ground state and the 19385, 19746, 19572, and 19575 cm^{−1} excited states, respectively. Strong variation of the effective Ω quantum number with J has been detected for the lowest rotational levels of the 19572 cm^{−1} state. The effective values have been determined to be equal to Ω = 1 (fixed value), Ω = 1.472(7), and Ω = 1.624(18) for J = 1, J = 2, and J = 3, respectively. Possible perturbers of the 19572 cm^{−1} state will be considered. Electronic character of the ground state and the four excited states will be discussed. Correlation of the dipole moment and the Ω quantum number in Stark spectroscopic measurements will be analyzed and required experimental conditions for breaking of such correlation will be determined.